The input voltage to a transformer is 24 V a.c. and the output voltage is 6.0 V. The input coil has 720 turns
How many turns are on the output coil?
1
Expert's answer
2020-12-17T07:38:19-0500

Answer

Relationship between input and output voltage and number of turns is given by

"\frac{V_p}{V_s}=\frac{N_p}{N_s}"

turns are on the output coil

"N_s=\frac{N_p V_s}{V_p}=\frac{720\times6}{24}=180"